OpenStudy (anonymous):

Find the sum. 2x/x^2-6x+9 (plus) 4/x^2+2x-15

OpenStudy (anonymous):

2x/ ((x-3)^2 + 4/((x+5)(x-3)) =(2x(x+5) + 4(x-3) ) / (x+5)(x-3)^2 =(2x^2 +10x +4x -12) / ((x+5)(x-3)^2) =(2x^2+14x-12) / ((x+5)(x-3)^2) =2(x^2 +7x-6) / ((x+5)(x-3)^2)

OpenStudy (anonymous):

Is that in simplest form?

OpenStudy (anonymous):

Yup , I think so. I do not think that it needs expansion for the denominator as it is not complex at all.